Problem 20-1A ThreePoint Sports Inc, manufactures basketballs for the Womens National Basketball Association (WNBA). For the(a) Prepare an incremental analysis for the special order. (Round all per unit computations to 2 decimal places, e.g. 15.25.
Answer #1
Reject order Accept order Net income increase(decrease)
Revenues 0 280000 280000
Cost of goods sold 0 -217600 -217600
Selling and administrative expenses 0 -23900 -23900
Net income 0 38500 38500
b
Yes accept the special order
c
Minimum selling price 29.23 =21.76+1.63+0.76+5.08
Workings:
Per unit:
Cost of goods sold 21.76 =(3569024-960000)/119900
Selling and administrative expenses 1.63 =(443437-248000)/119900
Revenues 280000 =10000*28
Cost of goods sold 217600 =10000*21.76
Selling and administrative expenses 23900 =10000*(1.63+0.76)